# Beyond Discretization: Learning the Optimal Solution Path

### v2

@@ A.3 Proofs from Section 4
As described in the main text, $\boldsymbol{\Phi}(\lambda)\in\mathbb{R}^{d\times qd}$ is block-diagonal with $d$ blocks of size $1\times q$. Denote this block by $\psi(\lambda)\in\mathbb{R}^{1\times q}$ and note the elements of $\psi(\lambda)$ are precisely the first $q$ Legendre polynomials evaluated at $\lambda$.
@@
Our proof is constructive and we will show a slightly stronger result. We will approximate $\boldsymbol{\theta}^{*}_{i}(\lambda)$ by the its Chebyshev truncation up to degree $q$ for each $i$. Letting $\bar{\boldsymbol{\beta}}$ be the coefficients corresponding to the resulting polynomials, we will show that $\epsilon_{\rm sp}(\bar{\boldsymbol{\beta}})$ satisfies the bound in Lemma 4.2, which implies that $\epsilon_{\rm sp}^{\ast}$ also satisfies the bound.
Let $T_{n}(\lambda)$ denote the $n^{\text{th}}$ Chebyshev polynomial of the first kind. For the $i^{\text{th}}$ dimension of the solution path, let $a_{i,n}$ denote the coefficient of $T_{n}$ in the Chebyshev truncation of $\boldsymbol{\theta}_{i}^{\ast}$ up to degree $q$.
Since $\boldsymbol{\theta}^{(\nu)*}_{i}(\lambda)$ is Lipschitz continuous with constant $V$, it also has bounded variation $V$. Thus, for any $q>\nu$ and $i\in[d]$, Trefethen (2013, Theorem 7.2) guarantees that
$\sup_{\lambda\in[-1,1]}\left|\sum_{n=0}^{q}a_{i,n}T_{n}(\lambda)-\boldsymbol{\theta}_{i}^{\ast}(\lambda)\right|~{}\leq~{}\frac{2V}{\pi\nu(q-\nu)^{\nu}}.$
@@
$\epsilon_{\rm sp}(\bar{\boldsymbol{\beta}})~{}=~{}\sup_{\lambda\in\Lambda}\left\{h\left(\left(\sum_{n=0}^{q}a_{i,n}T_{n}(\lambda)\right)_{i=[d]},\lambda\right)-h(\boldsymbol{\theta}^{*}(\lambda),\lambda)\right\}.$
@@
$h\left(\left(\sum_{n=0}^{q}a_{i,n}T_{n}(\lambda)\right)_{i=[d]},\lambda\right)-h(\boldsymbol{\theta}^{*}(\lambda),\lambda)~{}\leq~{}\frac{L}{2}\left\|\left(\sum_{n=0}^{q}a_{i,n}T_{n}(\lambda)\right)_{i=[d]}-\boldsymbol{\theta}^{*}(\lambda)\right\|^{2}.$
@@
$\epsilon_{\rm sp}(\bar{\boldsymbol{\beta}})$ $~{}\leq~{}\sup_{\lambda\in\Lambda}\left\{\frac{L}{2}\left\|\left(\sum_{n=0}^{q}a_{i,n}T_{n}(\lambda)\right)_{i=[d]}-\boldsymbol{\theta}^{*}(\lambda)\right\|^{2}\right\}$
$~{}\leq~{}\sup_{\lambda\in[-1,1]}\left\{\frac{dL}{2}\cdot\sup_{i\in[d]}\left|\sum_{n=0}^{q}a_{i,n}T_{n}(\lambda)-\boldsymbol{\theta}_{i}^{\ast}(\lambda)\right|^{2}\right\}$
@@
$\frac{8L(\bar{\eta}+1)}{\mu}\frac{C_{q}}{c_{q}}\cdot\epsilon_{\rm sp}^{\ast}(q)\lesssim q^{2(1-\nu)},$
for some constant $A_{1}$ not depending on $\epsilon$. Hence it suffices to take
@@
$\frac{L}{\bar{\eta}\mu}\frac{C_{q}}{c_{q}}\log\left(\frac{c_{q}\mu\|\boldsymbol{\beta}^{*}_{\rm avg}\|^{2}}{2\bar{\eta}\epsilon_{\rm sp}^{\ast}}\right)\ \lesssim\ q^{2}\left(\log\|\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}\|^{2}+\log(q)\right)$
where we have assumed SGD was initialized at $\boldsymbol{\beta}=0$ and used Equation 10 and $c_{q}=\frac{2}{2q-1}$ (cf. the proof of Lemma 4.1) to simplify.
To complete the proof we need to bound $\|\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}\|^{2}$. To this end, we first bound $\|\bar{\boldsymbol{\beta}}\|$, where $\bar{\boldsymbol{\beta}}$ is constructed by approximating each component $\theta_{i}^{*}(\lambda)$ by its Chebyshev truncation to degree $q$ for each $i$.
@@
$\|\bar{\boldsymbol{\beta}}\|^{2}$ $=\sum_{i=1}^{d}\sum_{k=1}^{q}\left|a_{ik}\right|^{2}$
@@
where the penultimate inequality collects constants and invokes Trefethen (2013, Theorem 7.1) to bound the second summation. Note, for $\nu\geq 0$, this last sum is summable, so that $\|\bar{\boldsymbol{\beta}}\|\lesssim 1$.
Furthermore, from the proof of Lemma 4.2, $\epsilon_{\rm sp}(\bar{\boldsymbol{\beta}})\lesssim q^{-2\nu}$. Hence, since the solution-path error upper bounds the stochastic error, we also have $F(\bar{\boldsymbol{\beta}})-F(\boldsymbol{\beta}_{\mathrm{avg}}^{\ast})\lesssim q^{-2\nu}$. From strong convexity, this implies that $\|\bar{\boldsymbol{\beta}}-\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}\|\ \lesssim\ q^{-\nu}$.
@@
$\sup_{\lambda\in[-1,1]}\left|\sum_{n=0}^{q}a_{i,n}T_{n}(\lambda)-\boldsymbol{\theta}_{i}^{\ast}(\lambda)\right|~{}\leq~{}\frac{2M\omega^{-q}}{\omega-1}.$
@@
$\epsilon_{\rm sp}(\bar{\boldsymbol{\beta}})~{}\leq~{}\sup_{\lambda\in[-1,1]}\left\{\frac{dL}{2}\cdot\sup_{i\in[d]}\left|\sum_{n=0}^{q}a_{i,n}T_{n}(\lambda)-\boldsymbol{\theta}_{i}^{\ast}(\lambda)\right|^{2}\right\}~{}\leq~{}\frac{dL}{2}\left(\frac{2M\omega^{-q}}{\omega-1}\right)^{2}.$
@@
Again, to bound the number of iterations, we will need to bound $\|\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}-\boldsymbol{\beta}_{0}\|$. We again will assume that $\boldsymbol{\beta}_{0}=\mathbf{0}$, and first consider bounding $\bar{\boldsymbol{\beta}}$. Recall, $\bar{\boldsymbol{\beta}}$ is obtained by approximating each component $i$ by the Chebyshev truncation of $\boldsymbol{\theta}^{*}(\lambda)$ to degree $q$. Then,
$\|\bar{\boldsymbol{\beta}}\|^{2}$ $=\sum_{i=1}^{d}\sum_{k=1}^{q}\left|a_{ik}\right|^{2}$
@@
Here, the second to last inequaity uses (Trefethen, 2013, Theorem 8.1). Since the last summation is summabble, we again conclude that $\|\bar{\boldsymbol{\beta}}\|_{2}\lesssim 1$.
The proof of Lemma 4.4 establishes that $\epsilon_{\rm sp}(\bar{\boldsymbol{\beta}})\lesssim\omega^{-2q}$. Since solution path error bounds the stochasstic error, we conclude that $F(\bar{\boldsymbol{\beta}})-F(\boldsymbol{\beta}_{\mathrm{avg}}^{\ast})\lesssim\omega^{-2q}$, and by strong convexity, $\|\bar{\boldsymbol{\beta}}-\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}\|\lesssim\omega^{-q}$.
@@
$q^{2}\left(\log\|\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}\|+\log q\right)$ $\lesssim\log^{2}(1/\bar{\epsilon})\log\log(1/\bar{\epsilon})$
@@ B.1 Calibrating Uniform Discretization
To apply uniform discretization in practice on must decide i) the number of grid points to use, and ii) the number of gradient calls to make at each grid point. Loosely speaking, our approach to these two decisions is to first fix a set of desired “target” solution path errors denoted $\Delta$. Then, for each $\delta\in\Delta$, we determine the number of grid points and gradient calls to approximately achieve a solution path error of $\delta$.
@@
$\sup_{\lambda\in G(\delta)}h(\hat{\boldsymbol{\theta}}(\lambda),\lambda)-h(\boldsymbol{\theta}^{*}(\lambda),\lambda).$
@@
$\epsilon_{\rm sp}(\hat{\boldsymbol{\theta}}(\cdot))=\sup_{\lambda\in\Lambda}h(\hat{\boldsymbol{\theta}}(\lambda),\lambda)-h(\boldsymbol{\theta}^{*}(\lambda),\lambda).$
@@
In Figure 4, we compare the solution path error and grid pass error to the target solution path error for our experiments. Figure 4(a) illustrates that in the weighted binary classification experiment, the discretization scheme performs well, satisfying both objectives as the solution path error and grid pass error closely align with the target solution path error. Figure 4(b) demonstrates that for the portfolio allocation problem, both solution path error and grid pass error are still close to each other, but fall below the target solution path error. This is primarily due to the high precision of L-BFGS, though only very few gradient calls are performed at each grid.
@@ Appendix C Additional Figures
Figure 5: $\boldsymbol{\theta}^{*}(\lambda)$ for Portfolio Allocation from Section 6.2. The true solution paths are very smooth in each dimension, suggesting that this problem is highly interpolable by a polynomial basis.
Figure 6: LSP and ALSP for Portfolio Allocation. Compares methods using the first $p=4,9,16$ bivariate-Legendre polynomials. Differs from Figure 3 as the $y$-axis records the number of function evaluations during line-search of L-BFGS instead of gradient calls.

### v3

@@ A.3 Proofs from Section 4
As described in the main text, $\bm{\Phi}(\lambda)\in\mathbb{R}^{d\times qd}$ is block-diagonal with $d$ blocks of size $1\times q$. Denote this block by $\psi(\lambda)\in\mathbb{R}^{1\times q}$ and note the elements of $\psi(\lambda)$ are precisely the first $q$ Legendre polynomials evaluated at $\lambda$.
@@
Our proof is constructive and we will show a slightly stronger result. We will approximate $\bm{\theta}^{*}_{i}(\lambda)$ by its Chebyshev truncation up to degree $q$ for each $i$. Letting $\bar{\bm{\beta}}$ be the coefficients corresponding to the resulting polynomials, we will show that $\epsilon_{\rm sp}(\bar{\bm{\beta}})$ satisfies the bound in Lemma 4.2, which implies that $\epsilon_{\rm sp}^{\ast}$ also satisfies the bound.
Let $T_{n}(\lambda)$ denote the $n^{\text{th}}$ Chebyshev polynomial of the first kind. For the $i^{\text{th}}$ dimension of the solution path, let $a_{i,n}$ denote the coefficient of $T_{n}$ in the Chebyshev truncation of $\bm{\theta}_{i}^{\ast}$ up to degree $q$.
We assumed that $\bm{\theta}^{*}_{i}(\lambda)$, …, $\bm{\theta}^{(\nu-1)*}_{i}(\lambda)$ are absolutely continuous and $\bm{\theta}^{(\nu)*}_{i}(\lambda)$ has bounded variation $V$. When $\nu=0$, we are assuming simply that $\bm{\theta}^{*}_{i}(\lambda)$ is of bounded variation $V$. Thus, for any $q\geq\nu+1$ and $i\in[d]$, Trefethen (2013, Theorem 7.2) guarantees that
$\sup_{\lambda\in[-1,1]}\left|\sum_{n=0}^{q}a_{i,n}T_{n}(\lambda)-\bm{\theta}_{i}^{\ast}(\lambda)\right|~{}\leq~{}\frac{2V}{\pi\nu(q-\nu)^{\nu}}.$
@@
$\epsilon_{\rm sp}(\bar{\bm{\beta}})~{}=~{}\sup_{\lambda\in\Lambda}\left\{h\left(\left(\sum_{n=0}^{q}a_{i,n}T_{n}(\lambda)\right)_{i=[d]},\lambda\right)-h(\bm{\theta}^{*}(\lambda),\lambda)\right\}.$
@@
$h\left(\left(\sum_{n=0}^{q}a_{i,n}T_{n}(\lambda)\right)_{i=[d]},\lambda\right)-h(\bm{\theta}^{*}(\lambda),\lambda)~{}\leq~{}\frac{L}{2}\left\|\left(\sum_{n=0}^{q}a_{i,n}T_{n}(\lambda)\right)_{i=[d]}-\bm{\theta}^{*}(\lambda)\right\|^{2}.$
@@
$\epsilon_{\rm sp}(\bar{\bm{\beta}})$ $~{}\leq~{}\sup_{\lambda\in\Lambda}\left\{\frac{L}{2}\left\|\left(\sum_{n=0}^{q}a_{i,n}T_{n}(\lambda)\right)_{i=[d]}-\bm{\theta}^{*}(\lambda)\right\|^{2}\right\}$
$~{}\leq~{}\sup_{\lambda\in[-1,1]}\left\{\frac{dL}{2}\cdot\sup_{i\in[d]}\left|\sum_{n=0}^{q}a_{i,n}T_{n}(\lambda)-\bm{\theta}_{i}^{\ast}(\lambda)\right|^{2}\right\}$
@@
We note that this result is presented and holds for any $\nu\geq 0$. (When $\nu=0$, the bound is infinity and hence trivially valid.) As will be seen in the proof of Theorem 4.3 below, however, we will only apply the result when $\nu\geq 2$.
@@
$\frac{8L(\bar{\eta}+1)}{\mu}\frac{C_{q}}{c_{q}}\cdot\epsilon_{\rm sp}^{\ast}(q)\lesssim q^{2(1-\nu)}.$
Hence it suffices to take
@@
$\frac{L}{\bar{\eta}\mu}\frac{C_{q}}{c_{q}}\log\left(\frac{c_{q}\mu\|\bm{\beta}^{*}_{\rm avg}\|^{2}}{2\bar{\eta}\epsilon_{\rm sp}^{\ast}}\right)\ \lesssim\ q^{2}\left(\log\|\bm{\beta}_{\mathrm{avg}}^{\ast}\|^{2}+\log(q)\right)$
where we have assumed SGD was initialized at $\bm{\beta}=0$ and used Equation 10 and $c_{q}=\frac{2}{2q-1}$ (cf. the proof of Lemma 4.1) to simplify.
To complete the proof we need to bound $\|\bm{\beta}_{\mathrm{avg}}^{\ast}\|^{2}$. To this end, we first bound $\|\bar{\bm{\beta}}\|$, where $\bar{\bm{\beta}}$ is constructed by approximating each component $\theta_{i}^{*}(\lambda)$ by its Chebyshev truncation to degree $q$ for each $i$.
@@
$\|\bar{\bm{\beta}}\|^{2}$ $=\sum_{i=1}^{d}\sum_{k=1}^{q}\left|a_{ik}\right|^{2}$
@@
where the penultimate inequality collects constants and invokes Trefethen (2013, Theorem 7.1) to bound the second summation. Note, for $\nu\geq 0$, this last sum is summable, so that $\|\bar{\bm{\beta}}\|\lesssim 1$.
Furthermore, from the proof of Lemma 4.2, $\epsilon_{\rm sp}(\bar{\bm{\beta}})\lesssim q^{-2\nu}$. Hence, since the solution-path error upper bounds the stochastic error, we also have $F(\bar{\bm{\beta}})-F(\bm{\beta}_{\mathrm{avg}}^{\ast})\lesssim q^{-2\nu}$. From strong convexity, this implies that $\|\bar{\bm{\beta}}-\bm{\beta}_{\mathrm{avg}}^{\ast}\|\ \lesssim\ q^{-\nu}$.
@@
$\sup_{\lambda\in[-1,1]}\left|\sum_{n=0}^{q}a_{i,n}T_{n}(\lambda)-\bm{\theta}_{i}^{\ast}(\lambda)\right|~{}\leq~{}\frac{2M\omega^{-q}}{\omega-1}.$
@@
$\epsilon_{\rm sp}(\bar{\bm{\beta}})~{}\leq~{}\sup_{\lambda\in[-1,1]}\left\{\frac{dL}{2}\cdot\sup_{i\in[d]}\left|\sum_{n=0}^{q}a_{i,n}T_{n}(\lambda)-\bm{\theta}_{i}^{\ast}(\lambda)\right|^{2}\right\}~{}\leq~{}\frac{dL}{2}\left(\frac{2M\omega^{-q}}{\omega-1}\right)^{2}.$
@@
Again, to bound the number of iterations, we will need to bound $\|\bm{\beta}_{\mathrm{avg}}^{\ast}-\bm{\beta}_{0}\|$. We again will assume that $\bm{\beta}_{0}=\mathbf{0}$, and first consider bounding $\bar{\bm{\beta}}$. Recall, $\bar{\bm{\beta}}$ is obtained by approximating each component $i$ by the Chebyshev truncation of $\bm{\theta}^{*}(\lambda)$ to degree $q$. Then,
$\|\bar{\bm{\beta}}\|^{2}$ $=\sum_{i=1}^{d}\sum_{k=1}^{q}\left|a_{ik}\right|^{2}$
@@
Here, the second to last inequality uses (Trefethen, 2013, Theorem 8.1). Since the last summation is summable, we again conclude that $\|\bar{\bm{\beta}}\|_{2}\lesssim 1$.
The proof of Lemma 4.4 establishes that $\epsilon_{\rm sp}(\bar{\bm{\beta}})\lesssim\omega^{-2q}$. Since solution path error upper-bounds the stochastic error, we conclude that $F(\bar{\bm{\beta}})-F(\bm{\beta}_{\mathrm{avg}}^{\ast})\lesssim\omega^{-2q}$, and by strong convexity, $\|\bar{\bm{\beta}}-\bm{\beta}_{\mathrm{avg}}^{\ast}\|\lesssim\omega^{-q}$.
@@
$q^{2}\left(\log\|\bm{\beta}_{\mathrm{avg}}^{\ast}\|+\log q\right)$ $\lesssim\log^{2}(1/\bar{\epsilon})\log\log(1/\bar{\epsilon})$
@@ B.1 Calibrating Uniform Discretization
To apply uniform discretization in practice one must decide i) the number of grid points to use, and ii) the number of gradient calls to make at each grid point. Loosely speaking, our approach to these two decisions is to first fix a set of desired “target” solution path errors denoted $\Delta$. Then, for each $\delta\in\Delta$, we determine the number of grid points and gradient calls to approximately achieve a solution path error of $\delta$.
@@
$\sup_{\lambda\in G(\delta)}h(\hat{\bm{\theta}}(\lambda),\lambda)-h(\bm{\theta}^{*}(\lambda),\lambda).$
@@
$\epsilon_{\rm sp}(\hat{\bm{\theta}}(\cdot))=\sup_{\lambda\in\Lambda}h(\hat{\bm{\theta}}(\lambda),\lambda)-h(\bm{\theta}^{*}(\lambda),\lambda).$
@@
Figure 5 compares the solution path error and grid pass error to the target solution path error for our experiments. Figure 5(a) illustrates that in the weighted binary classification experiment, the discretization scheme performs well, satisfying both objectives as the solution path error and grid pass error closely align with the target solution path error. Figure 5(b) demonstrates that for the portfolio allocation problem, both solution path error and grid pass error are still close to each other, but fall below the target solution path error. This is primarily due to the high precision of L-BFGS, though only very few gradient calls are performed at each grid.
B.2 Basis for Moderate Dimensional Portfolio Allocation
Recall that this experiment uses the following objective function:
$h(\theta,\lambda)\ =\ -\lambda_{1}\cdot\mu^{\top}\theta+\lambda_{2}\cdot\theta^{\top}\Sigma\theta+\|\theta-\lambda_{3:12}\|_{2}^{2},$
where $\lambda_{3:12}$ represents the current holdings.
Since this is a quadratic objective, let us first directly compute the optimal solution path
$\bm{\theta}^{*}(\lambda)\ =\ \frac{1}{2}\left(\bm{I}+\lambda_{2}\Sigma\right)^{-1}\left(\lambda_{1}\mu+2\lambda_{3:12}\right).$
Consider the eigendecomposition of $\Sigma$, which takes the form $\Sigma=P^{\top}DP$ for some orthonormal matrix $P$ and diagonal matrix $D=\text{diag}(d_{1},\ldots,d_{10})$. Plugging into the above closed form optimal solution path yields
$\bm{\theta}^{*}(\lambda)\ =\ \frac{1}{2}P^{\top}\text{diag}\left(\frac{1}{1+\lambda_{2}d_{1}},\ldots,\frac{1}{1+\lambda_{2}d_{10}}\right)P\left(\lambda_{1}\mu+2\lambda_{3:12}\right).$
Using the Taylor series expansion for $\frac{1}{1+a\lambda_{2}}$ around 00
$\frac{1}{1+a\lambda_{2}}=\sum_{i=0}^{\infty}(-1)^{i}(a\lambda_{2})^{i},\qquad a\in\mathbb{R},$
we conclude that, as $q\to\infty$, each component of the optimal solution path $\bm{\theta}^{*}_{i}(\lambda)$ is in the span of our chosen basis:
$\lambda_{(j\bmod 12)}\cdot\lambda_{2}^{\lfloor{j/12}\rfloor},\qquad j=1,...,q.$
@@ Appendix C Additional Figures
Figure 6: $\bm{\theta}^{*}(\lambda)$ for Portfolio Allocation from Section 6.2. The true solution paths are very smooth in each dimension, suggesting that this problem is highly interpolable by a polynomial basis.
Figure 7: LSP and ALSP for Portfolio Allocation. Compares methods using the first $q=4,9,16$ bivariate-Legendre polynomials. Differs from Figure 3 as the $y$-axis records the number of function evaluations during line-search of L-BFGS instead of gradient calls.
Figure 8: Probability Density Function(PDF) of Beta Distribution over $[0,1]$ with Various $a$, $b$.